### v2

# Altermagnetism-induced non-collinear superconducting diode effect and unidirectional superconducting transport

Abstract
Current studies of non-reciprocal superconducting (SC) transport have centered on the forward–backward asymmetry of the critical current measured along a single axis. In most realizations, this diode effect is achieved via introducing ferromagnetism or applying an external magnetic field, which drives system into an effective Fulde–Ferrell (FF) state but often at the cost of severely suppressing the SC gap and thus compromising device robustness. Here we propose and theoretically demonstrate that coupling a conventional $s$-wave SC thin film to a $d$-wave altermagnet offers a more resilient alternative. The momentum-dependent spin splitting inherent to altermagnets induces a non-collinear SC-diode effect in the BCS state, with the critical-current anisotropy exhibiting a fourfold ($C_{4}$) symmetry. Upon entering the FF state at large splitting, this anisotropy gradually evolves into a unidirectional ($C_{1}$) pattern. Crucially, the FF pairing momentum locks to the discrete crystal axes, eliminating the rotational Goldstone mode and preserving a sizable SC gap without any abrupt or significant suppression. These combined features make the altermagnetic proximity an appealing platform to engineer symmetry-protected, energy-efficient and programmable SC diodes for next-generation electronic devices.
Introduction.—Non-reciprocal superconducting (SC) transport, in which the supercurrent flows without dissipation in one direction but encounters finite resistance in the reverse, has attracted considerable and growing interest [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] for its potential applications in next-generation quantum and ultra-low-dissipation electronic technologies. This phenomenon, known as the SC diode effect, enables directional control of supercurrents and offers a promising pathway toward inherently programmable SC circuits. The underlying microscopic mechanism so far is associated with the finite-momentum Cooper pairing [4]. Realizing such a state typically requires breaking time-reversal symmetry, for instance, by introducing ferromagnetism or applying an external magnetic field. The resulting Zeeman spin splitting leads to mismatched Fermi surfaces for spin-up and spin-down electrons [Fig. 1(a)], favoring SC pairs with a finite center-of-mass (CM) momentum and thus driving a Fulde–Ferrell–Larkin–Ovchinnikov (FFLO) transition [15, 16]. As a consequence, the critical current for superconductivity along the pairing CM momentum differs from that in the opposite direction, giving rise to the diode effect.
However, several challenges hinder the practical implementation of this mechanism. The external magnetic fields typically employed to induce Zeeman splitting often introduce the additional orbital effect, and hence suppress superconductivity by generating vortices or pair-breaking currents [17]. On the other hand, in the original FFLO theory, the direction of the Cooper-pair CM momentum is energetically degenerate [16], such that the emergence of a CM momentum spontaneously breaks the continuous rotational symmetry and gives rise to a gapless rotational Goldstone mode [18, 19]. This, in principle, renders the conventional FFLO state inherently fragile and difficult to stabilize. To address this issue, various theoretical proposals have introduced spin–orbit coupling (SOC) to lift the degeneracy and pin the CM momentum direction [20, 21, 22, 23, 24, 25, 26]. However, the first-order FFLO transition is usually accompanied by a substantial SC gap suppression [20, 16, 21, 27], weakening the condensate and undermining the robustness required for diode functionality. This intrinsic trade-off severely limits the practical viability of the FFLO-based SC diodes.
Very recently, a novel class of magnetic materials, known as altermagnets, has been theoretically proposed and possibly observed experimentally [28, 29, 30, 31, 32, 33, 34, 35]. Altermagnets feature oppositely spin-polarized sublattices related by crystalline rotation symmetries. While they break time-reversal symmetry like ferromagnets, they exhibit no net magnetization due to the sublattice compensation, similar to antiferromagnets [28, 29, 30]. A key characteristic of altermagnetism is an even-parity, momentum-dependent spin splitting that respects inversion symmetry and originates from symmetry-protected band degeneracies, typically exhibiting $d$-, $g$-, or $i$-wave form [28, 29], resulting in unconventional spin textures, as illustrated in Fig. 1(b) for a $d$-wave case. These distinct features position altermagnets as highly promising platforms for both spintronic [36, 37, 38, 39, 40] and quantum transport applications [41, 42, 43, 44, 45].
Figure 1: Schematic of spin-up and spin-down Fermi surfaces under ferromagnetic (left) and $d$-wave altermagnetic (right) spin splitting.
Given their time-reversal-breaking yet inversion-symmetric character, and the absence of the orbital effects due to a net magnetization, altermagnets provide a natural platform for exploring their interplay with superconductivity [46, 47, 48, 49, 50, 51]. The intrinsic momentum-dependent spin polarization in altermagnets induces a mismatch between spin-up and spin-down Fermi surfaces, favoring the finite-momentum Cooper pairing even in the absence of external fields [52, 53]. Recent theoretical works [54, 55, 56, 57, 58] have investigated the stability of such pairing near the SC transition, typically by solving the linearized gap equation. A fully self-consistent microscopic analysis was performed by Hong et al. [27], who systematically compared conventional BCS pairing, chiral $p$-wave states, and FF pairing under comparable altermagnetic conditions. Their results revealed that, within a narrow but realistic range of altermagnetic strength, the FF state can become energetically favorable without requiring external magnetic fields or fine-tuned SOC.
Motivated by these developments [53, 52, 27], here we investigate directional SC transport in conventional $s$-wave SC thin films proximity-coupled [59, 60, 61, 62, 63] to a $d$-wave altermagnet, using a fully microscopic and self-consistent framework. As the altermagnetic strength increases from zero, we find that the critical current in the BCS regime evolves from being isotropic to strongly anisotropic with a fourfold ($C_{4}$) symmetry, reflecting the underlying crystalline symmetry of the altermagnetic spin texture. Thus, the directions of maximal and minimal critical currents are not antiparallel, indicating that the non-reciprocity deviates from the simple single-axis diode behavior and instead reflects a non-collinear SC diode effect.
Upon further increase of the altermagnetic strength, the system undergoes a transition into the FF state, where Cooper pairs acquire a finite CM momentum. In this regime, the critical-current contour gradually develops into a pronounced unidirectional ($C_{1}$) form aligned with the CM momentum, i.e., a single SC transport direction is spontaneously selected and pinned. Notably, unlike the conventional FFLO mechanism [20, 16, 21], where the emerging CM momentum from the Zeeman-split Fermi surfaces spontaneously breaks the continuous rotational symmetry, the momentum-dependent spin splitting in altermagnets is intrinsically encoded in the crystal lattice. As a result, the optimal pairing momentum is locked to a discrete set of directions (four in the $d$-wave case), eliminating the associated rotational Goldstone mode [18, 19] and stabilizing a directionally pinned SC state. Importantly, the BCS-to-FF transition driven by the altermagnetic splitting does not exhibit the abrupt and strong gap suppression typically observed in the conventional FFLO phases [20, 16, 21]. These features make altermagnets a promising platform for realizing robust, directionally controllable SC diode effects while maintaining a sizable SC gap and strong superfluid density, ensuring the operational stability and reliable integration required for high-performance, scalable SC electronics.
Model.—We begin with a conventional $s$-wave SC Hamiltonian, incorporating the $d$-wave altermagnetic spin splitting induced via proximity coupling [59, 60, 61] to a $d$-wave altermagnet (such as RuO2[35], ${\mathrm{Mn}}_{5}$${\mathrm{Si}}_{3}$ [31] and $\beta$-MnO2 [29, 64]). The Hamiltonian is written as [17, 65, 66, 59]
$H=\!\sum_{{\bf k}s}(\xi_{\bf k}+sh_{\bf k}){\hat{n}}_{{\bf k}s}+\!\sum_{\bf kk^{\prime}k_{1}k_{1}^{\prime}}U_{\bf kk^{\prime}k_{1}k_{1}^{\prime}}c_{{\bf k}\uparrow}^{\dagger}c_{{\bf k^{\prime}}\downarrow}^{\dagger}c_{{\bf k_{1}}\downarrow}c_{{\bf k_{1}^{\prime}}\uparrow},$ (1)
where number operator ${\hat{n}}_{{\bf k}s}=c_{{\bf k}s}^{\dagger}c_{{\bf k}s}$ with $c_{{\bf k}s}^{\dagger}$ and $c_{{\bf k}s}$ being the electron creation and annihilation operators for momentum $\bf k$} and spin $s=\pm$ ($\uparrow$ and $\downarrow$); $\xi_{\bf k}={\hbar^{2}k^{2}}/({2m})-\mu$ is the kinetic energy measured from the chemical potential $\mu$, with $m$ being the effective mass of electrons; $h_{\bf k}=\gamma_{d}(k_{x}^{2}-k_{y}^{2})$ denotes the $d$-wave momentum-dependent spin splitting [40, 28, 29, 35, 48] arising from the $d$-wave altermagnetic proximity effect. The second term describes a generalized spin-conserving interaction, which, for an on-site $s$-wave pairing interaction [67], takes the form $U_{\bf kk^{\prime}k_{1}k_{1}^{\prime}}=U\delta({\bf k+k^{\prime}-k_{1}-k_{1}^{\prime}})$.
We consider the phase-modulated FF state [20, 16, 21, 27, 68], characterized by a SC order parameter with a single finite CM momentum $\bf q$}, i.e., all Cooper pairs formed between two electrons with opposite spins and momenta $\bf k$} and $\bf k$$^{\prime}$ condense at the same finite CM momentum $\bf q$}, in order to maximize the available pairing phase space. Unlike pair-density-wave states (such as the multiple-$\bf q$} LO state [15]), which exhibit a spatial modulation of the gap and break translational symmetry, the FF state preserves the translational invariance of the SC gap, featuring only a spatially varying SC phase. The corresponding SC gap is given by $\Delta_{\bf q}=U\sum_{\bf k}\langle{c_{{\bf-k+q}\downarrow}c_{{\bf k+q}\uparrow}}\rangle$, which leads to the mean-field Hamiltonian:
$H_{\bf q}$ $=$ $\sum_{\bf k}\big[(\xi_{\bf k+q}+h_{\bf k+q}){\hat{n}}_{{\bf k+q}\uparrow}+(\xi_{\bf-k+q}-h_{\bf-k+q}){\hat{n}}_{{\bf-k+q}\downarrow}$ (2)
$+\Delta_{\bf q}c_{{\bf-k+q}\downarrow}c_{{\bf k+q}\uparrow}+\Delta_{\bf q}c_{{\bf k+q}\uparrow}^{\dagger}c_{{\bf-k+q}\downarrow}^{\dagger}\big]-\frac{\Delta_{\bf q}^{2}}{U}.$
Diagonalizing this Hamiltonian via Bogoliubov transformation, one obtains $H_{\bf q}=\sum_{\bf k}(E^{+}_{\bf k}\alpha_{\bf k}^{\dagger}\alpha_{\bf k}-E_{\bf k}^{-}\beta_{\bf k}^{\dagger}\beta_{\bf k}-E_{\bf k}2V_{\bf k}^{2})-\frac{\Delta_{\bf q}^{2}}{U}$. The SC free energy of the system is given by [69]
$\mathcal{F}^{s}_{\mathbf{q}}$ $=$ $\langle{H}\rangle+k_{B}T\sum_{{\bf k},\lambda=\pm,\eta=\pm}\big[f(\eta\lambda{E_{\bf k}^{\lambda}})\ln{f(\eta\lambda{E_{\bf k}^{\lambda}})}\big]$ (3)
$=$ $\sum_{{\bf k},\lambda=\pm}\big[(E_{\bf k}+\lambda\Gamma_{\bf k})f(E_{\bf k}+\lambda\Gamma_{\bf k})+k_{B}T{\ln}f(E_{\bf k}+\lambda\Gamma_{\bf k})$
$+(E_{\bf k}+\lambda\Gamma_{\bf k})\big(1-f(E_{\bf k}+\lambda\Gamma_{\bf k})\big)\big]-\sum_{\bf k}E_{\bf k}2v_{\bf k}^{2}-\frac{\Delta_{\bf q}^{2}}{U}$
$=$ $-\!\frac{1}{\beta}\sum_{\mathbf{k},\lambda=\pm}\ln\big(1\!+\!e^{-\beta\lambda E_{\mathbf{k}}^{\lambda}}\big)\!-\!\sum_{\mathbf{k}}\left(E_{\bf k}\!-\!\Xi_{\mathbf{k}}\right)\!-\!\frac{|\Delta_{\mathbf{q}}|^{2}}{U},$
where $\beta=1/(k_{B}T)$. Minimizing $\mathcal{F}^{s}_{\bf q}$ with respect to $\Delta_{\bf q}$ yields the self-consistent gap equation:
$\frac{1}{U}=\sum_{\bf k}\frac{f(E_{\bf k}^{+})-f(E^{-}_{\bf k})}{2E_{\bf k}},$ (4)
where the quasiparticle energy spectrum is given by $E_{\bf k}^{\pm}=\hbar{\bf v}_{\bf k}\cdot{\bf q}+h_{\bf k}\pm{E_{\bf k}}$ with ${\bf v}_{\bf k}=\hbar{\bf k}/m$ and $E_{\bf k}=\sqrt{\Xi_{k}^{2}+\Delta_{\bf q}^{2}}$ [27, 53]. Here, $\Xi_{\bf k}=\xi_{\bf k}+2\gamma_{d}(k_{x}q_{x}-k_{y}q_{y})$, suggesting a moving-frame transformation [70], while $f(x)$ is the Fermi-Dirac distribution function, and the condensate density $2{V_{\bf k}}^{2}=1-\Xi_{\bf k}/E_{\bf k}$. The quasiparticle energies here are tilted by the finite CM momentum. As a result, at $T=0$, the integrand in the gap equation becomes nonzero only when $E_{\bf k}^{+}>0$ and $E_{\bf k}^{-}<0$. Thus, SC anomalous pairing correlations vanish in momentum-space regions where either $E_{\bf k}^{+}<0$ or $E_{\bf k}^{-}>0$, as $f(E_{\mathbf{k}}^{+})-f(E_{\mathbf{k}}^{-})\equiv 0$. These regions are referred to as unpairing regions [16, 20, 71], where electrons no longer contribute to Cooper pairing and instead behave as normal-state excitations.
The equilibrium SC state is determined by self-consistently solving the gap equation and minimizing the free energy $\mathcal{F}^{s}_{\bf q}$ with respect to the pairing momentum $\bf q$}. This procedure allows us to determine the optimal CM momentum ${\bf q}_{0}$ and the corresponding practical SC gap $\Delta=\Delta_{\bf q=q_{0}}$ as a function of system parameters such as the altermagnetic strength and temperature. To investigate the transport properties, we introduce an additional SC momentum ${\bf p}_{s}=\nabla_{\bf R}\theta/2-e{\bf A}$ [72, 73, 74, 75, 71, 76, 65, 17, 77, 78], corresponding to a uniform phase gradient, superflow, or external vector potential. The resulting supercurrent is given by ${\bf j}_{s}=(e/m)n_{s}{\bf p}_{s}$ [72, 73, 74, 75, 71, 76, 65, 79], where $n_{s}$ is the superfluid density [80], which approaches the total electron density $n$ in the low-temperature limit [76, 68, 20, 72, 17, 81, 82]. To determine the critical current, we gradually increase the magnitude of ${\bf p}_{s}$ along a fixed direction and solve the gap equation self-consistently under the total momentum ${\bf q}_{0}+{\bf p}_{s}$. The SC gap collapses at a critical value ${\bf p}_{s,c}$, signaling the pair-breaking threshold and marking the onset of dissipation.
For comparison, we also consider a conventional FFLO model with an isotropic spin splitting arising from ferromagnetic proximity, modeled as $h_{\bf k}=\gamma_{s}k^{2}$. This serves to isolate the role of the anisotropic, crystal-symmetry-protected spin splitting characteristic of altermagnets.
Figure 2: Simulated order-parameter (OP) results for the optimal CM momentum $q_{0}$ and the corresponding practical SC gap $\Delta=\Delta_{\bf q=q_{0}}$ as functions of the spin-splitting strength for (a) ferromagnetic and (b) $d$-wave altermagnetic proximity. Insets: the $\Delta_{\bf q}$ landscape plotted in $q_{x}$$q_{y}$ plane, which reflects the inverse of the free-energy profile. The plotted $q$ range extends up to $0.8q_{0}$ in each direction. T=0.1T=0.1~K. Normalized parameters used here are: the zero-splitting gap $\Delta_{0}=2$~meV, coherence length $\xi_{0}=v_{F}/\Delta_{0}$, and splitting unit $\gamma_{0}=\Delta_{0}/k_{F}^{2}$.
Phase diagram.— As a specific example, we consider a typical SC thin film, NbN, which has a SC gap $\Delta_{0}\sim 2~\text{meV}$ at $T\sim 0$ and a critical SC temperature $T_{c}\sim 12$-15 K [68, 83, 84]. Due to the Cooper instability of interacting fermions underlying the BCS theory [17, 66, 65, 85], the momentum summation for the pairing electrons is restricted such that the spin-up electron with momentum $\bf k+q$} and the spin-down electron with momentum $\bf-k+q$} both lie within their respective Debye momentum shells around the Fermi surface, enforced by the spin-dependent condition $|\xi_{s{\bf k+q}}+sh_{s{\bf k+q}}|\leq\omega_{D}$. The CM momentum dependence of $\Delta_{\bf q}$ directly reflects the inverse of the free-energy landscape since $\mathcal{F}^{s}_{\mathbf{q}}-\mathcal{F}^{n}_{\mathbf{q}}\propto-\Delta_{\bf q}^{2}$, offering an efficient means for following analysis.
The results for the ferromagnetic-proximity case are shown in Fig. 2(a). As the Zeeman splitting $\gamma_{s}$ is raised from zero, $\Delta=\Delta_{\bf q=0}$ in the BCS state is only mildly suppressed due to the gradual reduction of the available pairing phase space. With further increase, as shown in the insets of Fig. 2(a), the $\Delta_{\bf q}$ landscape (the region with nonzero $\Delta_{\bf q}$) evolves from a single maximum at $q=0$ (BCS state), to a double-well structure with local maxima at both $q=0$ and $q\neq 0$, and eventually to a single maxima at a finite $q\neq 0$, signaling the formation of the FF state for $\gamma_{s}\geq 0.45\gamma_{0}$. This evolution reflects a first-order phase transition [86], and is accompanied by an abrupt and significant suppression of the SC gap $\Delta=\Delta_{\bf q=q_{0}}$. Owing to the energy degeneracy in the direction of ${\bf q}_{0}$, the FF state with an emerging CM momentum spontaneously breaks the continuous rotational symmetry and supports a gapless Goldstone mode [19, 18], rendering it intrinsically fragile to fluctuations and disorder. At higher splitting, the remaining $\Delta_{\bf q}$ landscape continuously shrinks, and the system undergoes a second-order transition to the normal state at $\gamma_{s}=0.55\gamma_{0}$. These results agree fully with the conventional FFLO theory [20, 21, 16].
The altermagnetic-proximity case exhibits qualitatively distinct behaviors, as shown in Fig. 2(b). As the $d$-wave spin splitting $\gamma_{d}$ increases from zero, the SC gap $\Delta=\Delta_{\bf q=0}$ in the BCS state is only mildly suppressed. After a critical strength $\gamma_{d}>0.6\gamma_{0}$, a finite CM momentum $q_{0}$ develops gradually, signaling a continuous transition to the FF state. Remarkably, this transition is not accompanied by any abrupt or substantial suppression of the SC gap $\Delta=\Delta_{\bf q=q_{0}}$. Both the gap reduction and the growth of $q_{0}$ evolve smoothly throughout the entire FF phase, in sharp contrast to the isotropic ferromagnetic case [Fig. 2(a)]. Beyond a second threshold $\gamma_{d}=0.66\gamma_{0}$, the system undergoes a first-order transition directly from the FF state to the normal state, evidenced by a sudden collapse of the SC gap. This markedly different topology (a second-order transition to the FF state and a first-order transition to the normal state) totally arises from the strongly anisotropic $d$-wave spin splitting, which reconstructs the available pairing phase space in a fundamentally different way from the isotropic case.
As seen from the insets of Fig. 2(b), tracking $\Delta_{\bf q}$ landscape [the regions with nonzero $\Delta_{\bf q}$ including both yellow and purple ones] as a function of increasing $\gamma_{d}$, we observe the following evolution. At zero splitting, the gap landscape is perfectly isotropic and circular, with a global maximum at ${\bf q}=0$, reflecting the full rotational symmetry of the BCS state. As $\gamma_{d}$ increases but remains below the critical value $0.6\gamma_{0}$, the circular contour gradually deforms into a nearly $C_{4}$-symmetric square-like shape in $\bf q$}-space, but with pronounced inward indentations along the spin-splitting axes (i.e., the $x$ and $y$ directions). This distortion mirrors the fourfold crystalline anisotropy of the $d$-wave altermagnet, although the global maximum remains pinned at the origin ($q=0$). Beyond the critical strength $0.6\gamma_{0}$, the gap at ${\bf q}=0$ (i.e., the BCS state) becomes fully suppressed, forming an isotropic depression around the origin. The system is thus poised to relocate the gap maximum to a finite momentum and enter the FF state. Immediately beyond this point, i.e., in the early FF regime, the $\Delta_{\bf q}$ landscape spans a continuous connected region with a circular depression at the center and pronounced inward indentations along the spin-splitting axes at the outer boundary. Thus, although the global maximum of $\Delta_{\bf q}$ has already shifted off center, the landscape still retains partial connectivity. As $\gamma_{d}$ increases further, this region begins to weaken in the continuity, and four incipient lobes become visible along the nodal directions of the splitting, in order to achieve available pairing phase space for the $d$-wave-split Fermi surfaces in Fig. 1(b). For $\gamma_{d}>0.63\gamma_{0}$, the landscape fractures into four discrete maxima, each pinned to a nodal direction of the $d$-wave spin splitting, thereby locking the FF pairing momentum [27] and enhancing the robustness against fluctuations by eliminating rotational Goldstone modes [19, 18]. Eventually, once the spin-split Fermi surfaces cease to overlap, the four lobes contract and $\Delta_{\bf q}$ collapses, signaling a first-order transition to the normal state.
Figure 3: De-pairing critical current $J_{c}(\delta\phi)$ as a function of the current direction at various spin splittings. (a) and (b): results for ferromagnetic proximity at different strengths $\gamma_{s}/\gamma_{0}$ in BCS and FF regimes, respectively; (c) and (d): results for $d$-wave altermagnetic proximity at different strengths $\gamma_{d}/\gamma_{0}$ in BCS and FF regimes, respectively. The characteristic current scale $J_{0}=en_{s}/(m\xi_{0})$ and then $J_{c}/J_{0}=p_{s,c}\xi_{0}$. The angle $\delta\phi=\phi_{\bf p_{s}}-\pi/4$ and we set $\phi_{\bf q_{0}}=\pi/4$ throughout.
Critical currents.—After determining the phase diagram, we now present the key transport result of this study: the angular dependence of the de-pairing current $J_{c}(\delta\phi)$, plotted in Fig. 3. The $\delta\phi=\phi_{\bf p_{s}}-\pi/4$ denotes the angle between the applied supercurrent direction and $\pi/4$ axis, which corresponds to one of the nodal directions of the $d$-wave altermagnetic spin splitting. As mentioned above, after the FF transition, the emerging optimal CM momentum ${\bf q}_{0}$ in the ferromagnetic-proximity case is directionally degenerate, whereas in the altermagnetic-proximity case it becomes locked to one of the discrete crystalline directions at $(2n+1)\pi/4$ axes. For concreteness, we set $\phi_{\bf q_{0}}=\pi/4$ throughout.
In the BCS phase at the ferromagnetic-proximity case [Fig. 3(a)], $J_{c}(\delta\phi)$ is fully isotropic and decreases gradually with increasing $\gamma_{s}$, as expected. Upon entering the FF phase via a first-order transition [Fig. 3(b)], the critical-current contour becomes strongly anisotropic, and a clear forward–backward asymmetry emerges on the ${\bf q}_{0}$ axis, i.e., $J_{c}(\delta\phi=0){\neq}J_{c}(\delta\phi=\pm\pi)$, manifesting as a SC-diode effect in agreement with earlier prediction [7]. Interestingly, the largest critical currents do not occur precisely at $\delta\phi=\pi$, but instead appear in the range $\delta\phi\approx 0.75$$0.5\pi$, while the minimum lies at $\delta\phi=0$. This shift, which has been overlooked in previous studies, has a geometric origin arising from the circular symmetry of the $\Delta_{\bf q}$ shell in the FF state. Specifically, in the conventional FF state, the gap landscape $\Delta_{\bf q}$ [e.g., the region with a nonzero $\Delta_{\bf q}$ in the inset of Fig. 2(a) for $\gamma_{s}=0.45\gamma_{0}$] is nearly degenerate along a circular shell in $\bf q$}-space, with the optimal pairing momentum ${\bf q}_{0}$ located near its outer edge. As a result, a supercurrent applied parallel to ${\bf q}_{0}$ quickly drives the total momentum to the edge of this shell and triggers the gap to collapse, yielding the minimum $J_{c}$. A current slightly off the antiparallel direction probes more stable segments of the shell, thereby resulting in the maximum $J_{c}$.
For an altermagnetic exchange field, the evolution of the critical current exhibits qualitatively distinct behavior. As the splitting increases from zero, the BCS state [Fig. 3(c)] gradually loses its isotropy, and the angular profile of $J_{c}(\delta\phi)$ starts to develop the pronounced oscillations. With increasing $\gamma_{d}$, the critical current along the spin-splitting axes (i.e., the $x$ and $y$ directions: $\delta\phi=\pm\pi/4$ and $\pm 3\pi/4$) is progressively suppressed, while the critical current along the nodal directions of the splitting (i.e., the diagonal directions: $\delta\phi=0,\pi$ and $\pm\pi/2$) is enhanced. As a result, $J_{c}(\delta\phi)$ acquires a pronounced fourfold ($C_{4}$) anisotropy, directly reflecting the $d$-wave character of the altermagnetic spin texture and corresponding to the $C_{4}$-symmetric, square-like shape of the $\Delta_{\bf q}$ landscape below the critical value [see the insets of Fig. 2(b)]. Thus, the maxima of $J_{c}(\delta\phi)$ occur along the diagonal directions, while the minima lie along $x$ and $y$ axes. This nontrivial angular misalignment between the directions of maximal and minimal critical currents gives rise to a non-collinear SC diode effect, where the non-reciprocity of the supercurrent is not aligned with a single transport direction, but instead emerges from distinct crystal axes. This non-collinear SC diode effect is robust, as it originates from the crystal-symmetry-protected nature of the altermagnetic spin splitting, enabling intrinsically stable and symmetry-guided rectification of supercurrents.
With further increase of $\gamma_{d}$, following the onset of the FF phase, a finite CM momentum ${\bf q}_{0}$ develops along one of the nodal directions of the altermagnetic spin splitting, and directly breaks SC transport along the opposite direction. This asymmetry occurs because ${\bf q}_{0}$ resides near the inner boundary of the gap landscape $\Delta_{\bf q}$, making the system particularly sensitive to supercurrents applied in the ${\bf q}_{0}$ direction, which easily push the total momentum beyond the de-pairing threshold and trigger SC breakdown. As a result, only three residual peaks remain in the critical-current profile, aligned with the other nodal directions ($\delta\phi=0,\pm\pi/2$), since in this early FF regime, the $\Delta_{\bf q}$ landscape still spans a continuous connected region in $\bf q$} space [see the inset of Fig. 2(b)], although its global maximum has already shifted away from the origin to ${\bf q}_{0}$. Notably, as $\gamma_{d}$ increases further, the critical-current peaks perpendicular to ${\bf q}_{0}$ are gradually suppressed. Once the $\Delta_{\bf q}$ landscape fractures into four discrete maxima at $\gamma_{d}\geq 0.63\gamma_{0}$ [see the inset of Fig. 2(b)], the critical-current contour evolves into a broadly peaked unidirectional ($C_{1}$) pattern aligned with ${\bf q}_{0}$, i.e., a single preferred SC transport direction is spontaneously selected and pinned. This continuous symmetry reduction in the critical-current profile from $C_{4}$ to $C_{1}$ symmetry, marked by four, then three, and eventually a single peak, is a direct consequence of the smooth-evolution nature of the BCS-FF transition in the $d$-wave altermagnetic system.
It should be emphasized that the altermagnetism-induced FF state gives rise to a significant genuine and nearly ideal non-reciprocal SC transport behavior, manifested by a pronounced critical-current peak around the ${\bf q}_{0}$ direction and an entirely vanishing SC response in the reverse, i.e., $J_{c}(\delta\phi=0)\neq 0$ and $J_{c}(\delta\phi=\pm\pi)=0$. This non-reciprocity is far more prominent than that in conventional FFLO states driven by isotropic splitting [Fig. 3(b)]. In particular, deep in the FF regime, SC transport is sustained only in the vicinity of the ${\bf q}_{0}$ direction, while all other directions are suppressed, further amplifying the non-reciprocal characteristics and enhancing their relevance for practical implementation. Notably, the altermagnetism-induced FF state does not exhibit any abrupt or substantial SC gap suppression, ensuring the persistence of a robust condensate even in the highly non-reciprocal regime.
Figure 4: Temperature–splitting phase diagram of the (a) SC gap and (b) the optimal CM momentum. (c) Overall phase diagram, where the BCS, normal, and FF phases are indicated in blue, red, and green, respectively.
Discussion.—The present study demonstrates that altermagnetic proximity offers a resilient and field-free route to realizing robust SC diode effects. The combined features revealed, including a non-collinear SC diode effect at weak spin splitting (i.e., in the BCS regime), pronounced SC non-reciprocity at large splitting (i.e., in the FF regime), and unidirectional SC transport deep in the FF state, along with a robust and sizable SC gap throughout the entire SC phase, make altermagnetic heterostructures promising platforms for symmetry-protected, energy-efficient SC electronics for next-generation quantum and ultra-low-dissipation circuit technologies. Physically, the gap equation at $q=0$ is an even function of $h_{\bf k}$ and remains invariant under $C_{4}$. One therefore expects that the SC instability and the resulting critical currents are $C_{4}$-symmetric. After the system develops a finite pairing momentum $\mathbf{q}$, the spontaneously selected direction of $\mathbf{q}$ reduces the rotational symmetry from $C_{4}$ to $C_{1}$, giving rise to the diode behavior.
While we have focused on low-temperature results to emphasize device performance, the framework can be straightforwardly extended to finite temperatures up to the critical temperature $T_{c}$. A complete temperature–splitting phase diagram of the order parameters (the SC gap and the optimal CM momentum) is presented in Fig. 4. As shown in the figure, with increasing temperature, (I) for $\gamma_{d}\in(0,0.2\gamma_{0})$, the system undergoes a second-order transition from the BCS to the normal state, characterized by a continuous suppression of the SC gap; (II) for $\gamma_{d}\in(0.2\gamma_{0},0.5\gamma_{0})$, the transition becomes first-order, marked by a discontinuous collapse of the gap, indicative of phase competition or metastability; (III) for $\gamma_{d}\in(0.5\gamma_{0},0.66\gamma_{0})$, the system first undergoes a second-order transition from the BCS to the finite-momentum FF state, followed by a first-order transition from the FF to the normal state, suggesting enhanced sensitivity to thermal fluctuations in the finite-momentum condensate. A tri-critical Lifshitz point appears at approximately $T\approx 7$ K and $\gamma_{d}\approx 0.5\gamma_{0}$, where three distinct phases (BCS, FF, and normal) coexist. This point simultaneously marks (i) the onset of a finite pairing momentum, (ii) the change in the nature of the SC transition, and (iii) a topology change in the phase diagram. Such a Lifshitz point reflects a qualitative restructuring of the SC ground state, and may manifest experimentally as a nonanalyticity in thermodynamic quantities or anomalous transport behavior near the transition[87, 88]. Moreover, in Supplement Materials [89] (see also references [90, 91] therein), we have also derived the emergence of spin-triplet pairing correlations [59, 92, 20] which may lead to novel phenomena such as an anomalous paramagnetic Meissner effect, as introduced in Refs. [93, 63] and experimentally observed in Ref. [94], and cause the magnetoelectric effect [20, 95].
Acknowledgments.—This work was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award Number DE-SC0020145, as part of the Computational Materials Sciences Program. F.Y. and L.Q.C. also acknowledge the generous support of Donald W. Hamer Foundation through a Hamer Professorship at Penn State.